2 lenses in contact made of materials with dispersive powers in the ratio 2:1, behave as an achromatic convex lens of focal length 10cm. The individual focal length of the lenses are

Grade:

1 Answers

Eshan
5 years ago
Dear student,

Since the dispersive powers are in the ratio 2:1, so are the magnitudes of focal lengths.

Hence\dfrac{f_1}{f_2}=-2 \implies f_1=-2f_2

Also the focal length of combination is 10cm.
\dfrac{1}{10}=\dfrac{1}{f_1}+\dfrac{1}{f_2}=-\dfrac{1}{2f_2}+\dfrac{1}{f_2}

\implies f_2=5cm , f_1=-10cm